OpenStudy (anonymous):

How do you factor? im going into PAP algebra2 and doing summer homework and i guess missed the day in previous classes where they teach you how to factor.

OpenStudy (anonymous):

what are you trying to factor?

OpenStudy (anonymous):

4x^2-9=0

OpenStudy (anonymous):

alrighty- i am sorry for giving you false hope. i cannot help you here but if you google it (i know its a really cruddy answer) but if you google it you will find the answer. i am so sorry:(

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i dont know what youre thanking me for but you are welcome- i guess:)

hero (hero):

Difference of squares a^2 - b^2 = (a + b)(a - b) In this case a = 2x , b = 3 Thus: 4x^2 - 9 = (2x)^2 - 3^2 = (2x + 3)(2x - 3) = 0

hero (hero):

@Mukjong13